You roll two six-sided dice. What is the probability that the sum is a multiple of 2 or a multiple of 3? Round your answer to the nearest tenth of a percent if necessary
Answer by greenestamps(13334) 

What do you need help with on this?
There are 6*6=36 possible outcomes on the two dice. List them all out and find the ones that have sums that are not multiples of either 2 or 3.
Then form the probability fraction, convert to percent, and round as directed.
1+1 = 2 no good, multiple of 2
1+2 = 3 no good, multiple of 3
1+3 = 4 no
1+4 = 5 OKAY
1+5 = 6 no
1+6 = 7 OKAY
2+1 = 3 no
2+2 = 4 no
...
You can do this without our help....
|
|
|